OpenStudy (anonymous):

factor 2x2 + 3x - 14

OpenStudy (campbell_st):

try (2x + 7)(x - 2)

OpenStudy (jmark):

Given the equation 2x^2 + 3x - 14 = 0, factorize to get value of x 2x^2 -4x+7x – 14 = 0 take the common term. 2x(x – 2) + 7(x=2) = 0 (x-2) (2x+7) = 0 equate the terms to zero (x-2)=0 , (2x+7) = 0 x = 0 , 2x = -7 x =0, x= -7/2 Maths questions for class 8 ICSE - http://icse.edurite.com/icse-question-papers/maths-question-papers-for-class-8-icse.html